OpenStudy (anonymous):

let m=2 what is the value of 36 m- 16 a.20 b. 40 c. 56 d. 346

OpenStudy (kamibug):

\(\Huge\color{#9f07f4}{🎶~Welcome~to~Openstudy!~🎶}\) When a variable touches a number, like in the problem 36m, it means to multiply. So since m = 2 it means 36m = 36 * 2

OpenStudy (anonymous):

thanks

rebeccaxhawaii (rebeccaxhawaii):

36*2=72-16=56

rebeccaxhawaii (rebeccaxhawaii):

so c is your final answer

OpenStudy (anonymous):

cool

rebeccaxhawaii (rebeccaxhawaii):

Please click "best response" for the person that helped you the most, this just shows the appreciation for the time they took to assist. If you want further help with anymore questions, please fan me so in the long run you can see when I'm on. I'll be sure to take a look at your question to help :)